OpenStudy (anonymous):

If f(x) = -x^2-9x, find f(-4)

OpenStudy (anonymous):

f(-4)=(-4)^2-9*(-4)=16+36=52

OpenStudy (anonymous):

wouldnt it be 20 because i got 20

OpenStudy (anonymous):

Yes that's right, I hadn't seen the minus sign in front of your x^2

OpenStudy (anonymous):

yay i was right! mind helping me in other problems i've got an algebra test tommorow

OpenStudy (anonymous):

Sure (I've got a math test tomorrow too haha)

OpenStudy (anonymous):

really? wow! i'll post it in another post